# Without Paired Labeled Data: An End-to-End Self-Supervised Paradigm for UAV-View Geo-Localization

### v1

@@ V CONCLUSION
This paper investigates a highly valuable and challenging task namely cross-view geo-localization based on UAV-view, aiming to reduce reliance on costly paired matching data. To address the challenges posed by cross-view discrepancies, we propose a lightweight unsupervised framework for UAV-view geo-localization, incorporating dynamic hierarchical memory learning and information consistency evolution learning. This framework unifies intra-view and cross-view feature representations in a high-dimensional space, further revealing their intrinsic correlations and cross-view mapping patterns. Experiments conducted on three public benchmark datasets demonstrate that our method, even without relying on paired matching data, surpasses some fully supervised counterparts while maintaining a compact network design. Furthermore, when adapted to the fully supervised setting, our approach achieves performance comparable to state-of-the-art models, further validating its effectiveness. These results enhance the practical applicability of UAV-view geo-localization and provide new opportunities for real-world deployment.

### v2

@@ V CONCLUSION
This paper investigates a highly valuable and challenging task namely self-supervised UAV-view cross-view geo-localization, aiming to reduce the reliance on paired UAV–satellite images. To this end, we propose a self-supervised UVGL method, which employs a shallow backbone with only 9 layers for feature extraction. Pseudo labels are generated via clustering, and a dual-branch contrastive learning mechanism is introduced to enhance the capability in representing intra-view visual patterns. To further improve intra-view discriminability and consistency, we design a dynamic hierarchical memory learning module. Additionally, to mitigate the structural and semantic domain gap between views, we propose an information consistency evolution learning module, which facilitates semantic alignment by learning high-dimensional feature consistency across views. To improve the quality and robustness of pseudo labels, we introduce a bidirectional feature consistency-based Pseudo Label Enhancement strategy, which refines pseudo-supervision signals and boosts cross-view representation learning. As a result, the proposed method effectively unifies both intra-view and cross-view feature representations in the embedding space. Extensive experiments conducted on three public benchmarks University-1652, SUES-200, and DenseUAV, demonstrate that our method significantly outperforms existing self-supervised methods, even surpassing several fully supervised methods in certain scenarios, without relying on paired UAV–satellite images. These results validate the effectiveness and generalization capability of our method, highlighting its potential for real-world UVGL applications.